From GlueXWiki
Jump to: navigation, search

July 30, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC update (Mike, Simon, Beni)
    • CDC wave form
  2. FDC update (Lubomir, Luke)
  3. Electronics (Fernando, Chris, Nick)
  4. Engineering
    • gas system [8]
  5. Other


Participants: Curtis, Mike (CMU), Eugene, Luke, Beni, Dave, Chris, Nick, Simon, and Lubomir (JLab).

CDC update

- Beni: grounding with a thick cable of the HVBs doesn't change anything in the noise: a typical noise event is shown above. Chris and Nick tried both connecting through 50 Ohms or without 50 Ohms (by removing it on some of the boards). Still the idea to be tested is to connect the grounds on many cards together.

- Nick: One HV channel trips. For now Nick checked that it doesn't come from the CAEN, or HV cable. Next check is to swap the HVB cards.

FDC update

- Lubomir showed cathode resolution as a function of the charge (first plot). Normally, it gets better at higher charges. This is the case for the resolution in x (wire-reconstruction resolution) which changes from 300 down to 220 microns. The resolution in y is almost constant with the charge. This indicates, there's a contribution to it at a level of <180 microns. Part of it may come from the tracking, but significant part is coming also from another source. Lat two plots show the cathode resolution (residuals with an offset for better colors). One clearly see the different foils on the cathodes. Also one can see the channel-by-channel gain calibration (done by Vlad) may help to improve the resolution.

- Garfield simulations showed that screening from the wire creates the asymmetry in the signals induced on the two cathodes, depending on the side w.r.t. the wire at which the avalanche is created. This is demonstrated in the plots above.

- Spare package: we took data with a random (pulser) trigger with the standard DAQ prepared by Sergey. We see iron source signal an every 200 events, still good enough to have some statistics now over much bigger area. After taking some data we changed the gas from 60/40 to 40/60 Ar/CO2.


- When we decided to connect the CDC and FDC mixing tanks, wrong connection was used resulting in filling the FDC only with Ar and putting all of the CO2 to the CDC. Dave fixed that and now the two chambers are approaching slowly 50/50. All this was checked with the CO2 sensor - see the plot.

- Dave is working on the interlock that shuts off the MFCs at the input of the FDC and CDC in case of pressure problems. For now, after such event it can be cleared only manually.